# EvaGaussians: Event Stream Assisted Gaussian Splatting from Blurry Images

### v1

@@ 3.3 Loss Functions
The formation of blur reconstruction loss is the same as in the original 3D-GS [9], it differs in utilizing blurry images as supervision and jointly optimizing the 3D-GS parameters and the camera trajectories, thus facilitating an initial deblurring reconstruction of 3D-GS.
@@
Event Reconstruction Loss. Leveraging the microsecond-level temporal resolution provided by the event streams, we further adopt an event reconstruction loss to aid in fine-grained deblurring. Specifically, we uniformly divide the exposure time into $m=n-1$ intervals, each with a length of $\frac{\tau}{m}$. Subsequently, we integrate the recorded event stream along these time intervals using Eq. 2, resulting in $m$ event maps $\{\mathbf{E}_{i}\}_{i=0}^{m-1}$ to serve as the event supervision. During training, for the $j$-th blurry view, we convert the rendered image sequence $\{\mathbf{\widetilde{I}}_{i}\}_{i=0}^{n-1}$ on the camera motion trajectory into event maps $\{\mathbf{\widetilde{E}}_{i}\}_{i=0}^{m-1}$, using a differentiable event simulator [31; 7], and constrain the discrepancies between the simulated event maps and the ground truth event maps, expressed as:
$\mathcal{L}_{event}=\frac{1}{m}\sum_{i=0}^{m-1}\|\mathbf{E}_{i}-\mathbf{\widetilde{E}}_{i}\|_{1}.$ (8)
The final loss function is the combination of the blur reconstruction loss and the event reconstruction loss, defined as:
$\mathcal{L}=\lambda_{blur}\mathcal{L}_{blur}+\lambda_{event}\mathcal{L}_{event}.$ (9)
During training, we jointly optimize the 3D-GS parameters and the camera motion trajectories using the defined loss function, ultimately achieving high-quality 3D-GS reconstruction from the captured blurry images and event streams.
3.4 Implementation Details.
We implemented EvaGaussians based on the official code of 3D-GS [9]. Throughout the training process, we set $\lambda_{1}=0.2$, $\lambda_{blur}=1$ and $\lambda_{event}=5\times{10}^{-3}$ for the loss function, and used $n=9$ for the number of poses to be optimized during the exposure time. In implementing the event reconstruction loss, we configured the positive threshold as ${c}_{pos}=0.25$ and the negative threshold as ${c}_{neg}=0.25$ for synthetic scenes, and set ${c}_{pos}=0.197$ and ${c}_{neg}=0.241$ for real scenes. The training process spans 50,000 iterations, with an event reconstruction loss introduced after a 3,000-iteration warmup. We execute 1,000 iterations to initialize the 3D-GS, after which we omit the densification process to streamline and simplify the subsequent optimization. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
4 Experiments
4.1 Datasets.
Referring to [30; 3], we evaluate our method on both synthetic and real-world data. To ensure a comprehensive evaluation, we have contributed two novel datasets:
EvaGaussians-Blender Dataset. To evaluate the model’s ability to handle different scales of scenes, we construct a synthetic dataset covering a variety of scene scales, coupling with camera trajectories and event data. For large-scale scenes, we use Blender to design five diverse scenes, including city blocks and natural sceneries. For medium-scale scenes, we design three additional scenes using Blender, including classrooms, bedrooms, and cafes; we also redesign the camera trajectories of four scenes from DeblurNeRF [20]. For object-level scenes, we create six scenes based on the dataset from NeRF[23] and ${\textnormal{E}^{2}}$eRF[30]. We simulate motion blur by manually placing multi-view cameras, randomly adjusting camera poses, and performing linear interpolation between the original and perturbed positions for each view. The images are rendered from these interpolated poses and blended in linear RGB space to produce the final blurred images. The corresponding event streams are generated using ESIM[31] and V2E[7]. The finally produced large-scale and medium-scale scenes include 35 views of blurred images and the corresponding event data, while the object-level scenes include 100 views of blurred images.
Table 1: Quantitative comparisons of novel view synthesis on large-scale scenes. The results are the average of 5 scenes, each with 100 views. We highlight the best-performing results in red and the second-best results in orange.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
Table 2: Quantitative comparisons of novel view synthesis on medium-scale scenes. The results are the average of 7 scenes, each with 100 views.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
Table 3: Quantitative comparisons of novel view synthesis for object-level Scenes. The results are the average of 6 objects, each with 100 views.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
[…]
EvaGaussians-DAVIS Dataset. The real-world experiments are performed on scenes captured by the Color DAVIS346 event camera [38]. It has a resolution of 346×260 pixels and we configured the exposure time for RGB frames at 100 milliseconds. We manually recorded five scenes, including three at the object level and two indoor scenes. The finally produced dataset consists of 30 images per scene coupled with the recorded event streams, each displaying various blur and lighting conditions. Please refer to Appendix. A for more details.
4.2 Experiment Settings
Baselines. We firstly compare our method with NeRF[23] and 3D-GS[9] directly trained on the blurry images, referring to them as Blurry-NeRF (B-NeRF) and Blurry-3DGS (B-3DGS). Then, we compare our method against several deblurring rendering methods, including ${\textnormal{E}^{2}}$eRF[30], BAD-NeRF[41], BAD-Gaussians[48] (BAD-GS), and EvDeblurNeRF[3] (EDNeRF). Among these, the first two methods simulate motion blur and optimize camera trajectories without relying on event data, whereas the latter two are event-assisted methods. Additionally, we employ three image deblurring methods, namely UFP (a single-image deblurring method), EDI[28] (an event-based deblurring method) and EFNet[37] (a learnable event-based deblurring method), to process the input blurry images. Subsequently, we train vanilla 3D-GS using the deblurred images and name the resulting baselines UFP-GS, EDI-GS, EFN-GS.
Evaluation Metrics. We utilize a set of metrics to assess the performance of our model. For synthetic datasets, we employ three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [42], and VGG-based Learned Perceptual Image Patch Similarity (LPIPS) [46]. These metrics serve to quantify the similarity between predicted novel views and provided target novel views. For real-world datasets, since the sharp ground-truth images are unavailable, we utilize a series of No-Reference Image Quality Assessment (NR-IQA) metrics to to evaluate our method. The adopted metrics include BRISQUE [24], NIQE [25], PIQE [40], RankIQA [18], and MetaIQA [50], which allow for effective evaluation when lacking ground truth images.
Figure 2: Qualitative comparision on the synthetic dataset. We show both novel view synthesis (NVS) results and input view deblurring (DVS) results on the top two rows. It shows that our method achieves better performance in recovering the training blurry views as well as rendering novel views. More results are presented in Appendix. B.
4.3 Synthetic Data Experiments
We evaluate our approach across a variety of scenes, including large-scale scenes, medium-scale scenes, and object-level scenes. Quantitative assessments of novel view synthesis are shown in Table. 1, Table. 2, and Table. 3, respectively. The deblurring results of input views are detailed in the Appendix. B. It can be found that our method achieves substantial improvements in most of the metrics, especially in challenging large scenes. Specifically, both B-NeRF and B-3DGS produce blurry novel views since they are directly trained on blurred images. The image deblurring-based baselines, UFP-GS, EDI-GS and EFN-GS, also produced inferior results, because the image deblurring process potentially corrupts the 3D consistency of the training images. Notably, our approach outperforms BAD-Gaussians[48] and BAD-NeRF[41], due to their limited capability in modeling complex textures. In addition, our method also surpasses the event-assisted methods ${\textnormal{E}^{2}}$eRF[30] and EvDeblurNeRF[3] in producing high-quality novel views with intricate details, with better training and rendering efficiency. An extended analysis of all the baselines is provided in the Appendix. B.
The qualitative results are illustrated in Figure. 2. It is found that that although ${\textnormal{E}^{2}}$eRF[30] performs well in object-level scenes, it struggles in medium and large-scale scene modeling, producing significant blurring results. Additionally, BAD-Gaussians[48] falls short in regions with pronounced color and depth variations, and produces overly smooth background textures. Although EvDeblurNeRF[3] exhibits overall satisfactory performance, its complex network architecture prolongs the training time (about 7 hours per scene) and precludes real-time rendering. In comparison, our method overcomes the baselines in producing high-fidelity novel views, and significantly reducing training time as well as demonstrating substantial advantages in real-time application scenarios. More visualization results are provided in Appendix. B.
@@ 4.3 Synthetic Data Experiments
Table 4: Quantitative comparisons of novel view synthesis on the real-world dataset. The table reports the average performance across five scenes. The results demonstrate that our method surpasses previous state-of-the-art approaches and achieves better performance across all metrics.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF    Ours
@@
Figure 3: Qualitative results on the real-world dataset. It can be found that our method outperforms the baselines in synthesizing sharper novel views. More results are presented in Appendix. B.
4.4 Real-world Data Experiments
We present the quantitative results on the captured real-world data in Table. 4. It can be found that our method achieves superior performance compared to other approaches. Specifically, for NR-IQA metrics, we achieve improvements in BRISQUE, NIQE, PIQE, and RankIQA by 15.38%, 19.50%, 11.49%, and 22.83% respectively. We also achieve an increase in 19.38% in MetaIQA. The qualitative comparisons are illustrated in Figure. 3, which further demonstrate that our method is capable of reconstructing detailed textures, ultimately achieving higher-quality novel view synthesis. More results are presented in Appendix. B.
4.5 Ablation study
Figure 4: Ablation on number of poses in the camera trajectory.
Table 5: Quantitative ablation on $\mathcal{L}_{\text{event}}$ and $\mathcal{L}_{\text{blur}}$.
w/o $\mathcal{L}_{\text{event}}\&\mathcal{L}_{\text{blur}}$ w/ $\mathcal{L}_{\text{blur}}$ w/ $\mathcal{L}_{\text{event}}\&\mathcal{L}_{\text{blur}}$
@@ 4.5 Ablation study
Figure 5: Qualitative ablation on $\mathcal{L}_{\text{event}}$ and $\mathcal{L}_{\text{blur}}$.
@@
Numbers of Camera Poses. We conduct ablations to investigate the effect of the number of camera poses optimized in the exposure time. We select five large scenes from our synthetic dataset for evaluation. In the experiments, we vary the number of camera poses, denoted as $n$, from 5, 9, 13, and 17. The quantitative results of the novel view rendering are displayed in Figure. 5. It indicates that the results reach a bottleneck at 9 poses. Beyond this point, the improvements are limited and may potentially lead to local convergence issues. Based on these experiments, we choose $n=9$ camera poses to achieve a balance between rendering performance and training efficiency. Here, we also provide comparison with BAD-NeRF [41] and BAD-Gaussians [48]. The two methods typically use linear interpolation to obtain camera trajectory, while our camera trajectories are estimated from the decomposed latent images, which provides more accurate initialization and helps our method achieves better performance.
@@
Blur Reconstruction Loss and Event Reconstruction Loss. We conduct novel view synthesis experiments on five large-scale scenes from our synthetic dataset to validate the training losses. The quantitative results are reported in Table. 5, which indicate that the blurry reconstruction loss and event reconstruction loss significantly enhance novel view rendering quality. The qualitative ablation are shown in Figure. 5, where without the supervision of the proposed loss functions, the reconstruction results tends to produce oversmoothed results. With only the blur reconstruction loss, the produced results also lack high-frequency details. With the full losses, our method can produce high-fidelity novel views with intricate details.

### v2

@@ 3.3 Event-assisted Bundle Adjustment
The formulation of blur reconstruction loss is the same as in the original 3D-GS [13], it differs in utilizing blurry images as supervision and jointly optimizing the 3D-GS attributes and the camera trajectories, thus facilitating an initial deblurring reconstruction of 3D-GS.
@@
Event Reconstruction Loss. Leveraging the abundant high-frequency information offered by the event streams, we further adopt an event reconstruction loss to aid in 3D-GS optimization. Specifically, we uniformly divide the exposure time into $m=n-1$ intervals, each with a duration of $\frac{\tau}{m}$. Subsequently, we integrate the recorded event stream along these time intervals using Eq. 1, resulting in $m$ event maps $\{\mathbf{E}_{i}\}_{i=1}^{m}$ to serve as event level supervision. During training, for the $j$-th blurry view, we convert the rendered image sequence $\{\mathbf{\widetilde{I}}_{i}\}_{i=1}^{n}$ on the camera trajectory into event maps $\{\mathbf{\widetilde{E}}_{i}\}_{i=1}^{m}$, using a differentiable event simulator [31, 9], and constrain the discrepancies between the simulated event maps and the ground truth event maps, expressed as:
$\mathcal{L}_{event}=\frac{1}{m}\sum_{i=1}^{m}\|\mathbf{E}_{i}-\mathbf{\widetilde{E}}_{i}\|_{1}.$ (6)
The event reconstruction loss further aids in recovering a sharp 3D-GS with improved texture details.
3.4 Event-assisted Geometry Regularization
The blurry color images are captured only during the exposure time and are much sparser than the event stream. Relying on such low-quality image-level supervision may cause the 3D-GS to overfit on the training images, resulting in significant floaters and inferior geometry, which affects the quality of novel view synthesis. Leveraging Eq. 1 and Eq. 4, given the continuously recorded event streams $\mathbf{E}(t)$, we can derive continuous grayscale intensity images $\mathbf{G}(t)$ that are rich in geometric information and can function beyond the exposure time. Motivated by this, we further propose two event-assisted geometry regularization terms to aid in 3D-GS training.
Intensity Reconstruction Loss. As shown in Figure. 2(C.1), during training, we randomly sample continuous time $t$ between the interval of two adjacent blurry image, and derive the grayscale intensity image $\mathbf{G}(t)$ using Eq. 2. We then minimize the difference between it and the rendered intensity image from 3D-GS, expressed as:
$\mathcal{L}_{int}=(1-\lambda_{2})\cdot\|\mathbf{G}(t)-\mathbf{\widetilde{G}}(t)\|_{1}+\lambda_{2}\cdot\text{D-SSIM}(\mathbf{G}(t),\mathbf{\widetilde{G}}(t)),$ (7)
where $\mathbf{\widetilde{G}}(t)$ is converted from the colored render result.
Intensity-aware Depth Regularization Loss. As shown in Figure. 2(C.2), to further improve the geometry of 3D-GS, inspired by [8, 4], we adopt an intensity-aware depth regularization loss during training, defined as:
$\mathcal{L}_{depth}=$ $\frac{1}{N}\sum_{x,y}(~{}|\partial_{x}\mathbf{\widetilde{D}}_{xy}(t)|e^{-\beta|\partial_{x}\mathbf{G}_{xy}(t)|}$ (8)
$+|\partial_{y}\mathbf{\widetilde{D}}_{xy}(t)|e^{-\beta|\partial_{y}\mathbf{G}_{xy}(t)|}~{}),$
where $\mathbf{\widetilde{D}}(t)$ is the rendered depth map, $(x,y)$ denotes the pixel location, $N$ is the total number of pixels, and $\beta$ is set to $2$ in our experiments. The horizontal and vertical gradients are calculated by applying convolution operations with $5\times 5$ Sobel kernels [39]. This regularization is founded on the observation that depth transitions in an image often correspond to changes in intensity. Therefore, it ensures that the spatial variation of depth closely matches that of the intensity image, thereby reducing geometric artifacts at object boundaries.
@@ 3.4 Event-assisted Geometry Regularization
Table 1: Quantitative comparisons of novel view synthesis across large-scale, medium-scale, object-level, and real-world scenes. The table reports the average performance for each scale, demonstrating that our method consistently surpasses previous state-of-the-art approaches across all metrics. Best-performing results are highlighted in bold and second-best results in underline.
Scene Type Metric B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@
[…]
The total loss function is the combination of the above losses, defined as:
$\mathcal{L}_{total}=$ $\lambda_{blur}\mathcal{L}_{blur}+\lambda_{event}\mathcal{L}_{event}$ (9)
$+\lambda_{int}\mathcal{L}_{int}+\lambda_{depth}\mathcal{L}_{depth}.$
4 Experiments
4.1 Implementation Details
Progressive Training. We implemented EvaGaussians based on the official code of 3D-GS [13]. The training process spans 50,000 iterations, with an event reconstruction loss introduced after a 3,000-iteration warmup and we omit the densification process to streamline and simplify the subsequent optimization. Additionally, we adopt a coarse-to-fine training strategy, starting with rendering at a low resolution (0.3$\times$ downsampling in the early 30% iterations) and progressively increasing the size of the rendered views to full resolution. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
Hyperparameter Setting. During the training process, we set $\lambda_{1}=0.2$, $\lambda_{blur}=1.0$, $\lambda_{depth}=1.0e^{-2}$, $\lambda_{event}=5.0e^{-3}$ and $\lambda_{int}=1.0e^{-3}$ for the loss function, and used $n=9$ for the number of poses to be optimized during the exposure time. In implementing the loss $\mathcal{L}_{event}$, we configured the positive threshold as ${c}_{pos}=0.25$ and the negative threshold as ${c}_{neg}=0.25$ for synthetic scenes, and set ${c}_{pos}=0.197$ and ${c}_{neg}=0.241$ for real scenes.
4.2 Datasets
To facilitate a comprehensive evaluation, we introduce two novel datasets, with an overview provided below. Detailed information is presented in the supplementary.
EvaGaussians-Blender Dataset. We construct a synthetic dataset covering a variety of scene scales, coupling with diverse camera trajectories and event data. For large-scale scenes, we employ Blender to craft five distinct scenes, including city blocks and natural landscapes. For medium-scale scenes, we craft three scenes using Blender, and redesign the camera trajectories of four scenes from DeblurNeRF [25]. For object-level scenes, we create six scenes based on the NeRF-synthetic [26] dataset. We simulate motion blur by manually placing multi-view cameras, randomly adjusting camera poses, and performing linear interpolation between the original and perturbed positions for each view. The images are rendered from these interpolated poses and blended in RGB space to produce the final blurry images. The corresponding event streams are simulated using ESIM [31] and V2E [9]. The resulting large-scale and medium-scale scenes comprise 35 views of blurry images along with their corresponding event data, whereas the object-level scenes feature 100 views of blurry images.
Figure 3: Qualitative comparison on the synthetic and real dataset. We show the rendering novel views on the top section (a) and exhibit both novel view synthesis results and input view deblurring results on the bottom section (b). It shows that our method achieves better performance in recovering the training blurry views as well as rendering novel views. More results are presented in the supplementary.
EvaGaussians-DAVIS Dataset. We manually recorded five real-world scenes using the Color DAVIS346 event camera [38], which has a resolution of $346\times 260$ pixels and an exposure time of 100 milliseconds for the RGB frames. The dataset includes three object-level scenes and two indoor scenes. After processing, the final dataset consists of 30 images per scene, along with the recorded event streams, each showcasing various blur and lighting conditions.
4.3 Experiment Settings
Baselines. We compare our method with three types of baselines: 1) NeRF [26] and 3D-GS [13] directly trained on the blurry images, referring to as B-NeRF and B-3DGS. 2) Deblur rendering methods, including BAD-NeRF [41], BAD-GS [52], ${\textnormal{E}^{2}}$eRF [30], and EDNeRF [3]. Among these, the first two methods simulate motion blur and optimize camera trajectories without event stream, whereas the latter two are event-assisted methods without camera trajectory optimization. 3) Image deblur methods, including UFP [6] (single-image deblurring), EDI [29] (event-based deblurring), and EFNet [35] (learnable event-based deblurring). We process input blurry images with them and train the vanilla 3D-GS with pre-deblurred images. The resulting baselines are referred to as UFP-GS, EDI-GS, and EFN-GS.
Evaluation Metrics. For synthetic datasets, we employ the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [42], and VGG-based Learned Perceptual Image Patch Similarity (LPIPS) [50] to evaluate the similarity between rendered novel views and ground-truth novel views. For real-world datasets, since the sharp ground-truth images are unavailable, we utilize several No-Reference Image Quality Assessment (NR-IQA) metrics for evaluation, including BRISQUE [27], NIQE [28], PIQE [40], RankIQA [21], and MetaIQA [53], which allow for image evaluation when lacking ground truth images.
4.4 Synthetic Data Experiments
We evaluate our approach across a variety of scenes, including large-scale scenes, medium-scale scenes, and object-level scenes. Quantitative assessments of novel view synthesis are shown in the first three rows of Table. 1. The deblurring results of input views are detailed in the supplementary. It can be found that our method achieves substantial improvements in most of the metrics, especially in challenging large scenes. Specifically, both B-NeRF and B-3DGS produce blurry novel views since they are directly trained on blurred images. The image deblurring-based baselines, UFP-GS, EDI-GS and EFN-GS, also produced inferior results, because the image deblurring process potentially corrupts the 3D consistency of the training images. Notably, our approach outperforms BAD-GS [52] and BAD-NeRF [41], due to their limited capability in modeling complex textures. In addition, our method also surpasses the event-assisted methods ${\textnormal{E}^{2}}$eRF [30] and EDNeRF [3] in producing high-quality novel views with intricate details, with better training and rendering efficiency. An extended analysis of all the baselines is provided in the supplementary.
The qualitative results are illustrated in Figure. 3, where the first three rows of Figure. 3(a) shows novel view synthesis results, and Figure. 3(b) shows both novel view and deblurring view synthesis results. More visualization results are provided in the supplementary. It can be found that although ${\textnormal{E}^{2}}$eRF [30] performs well in object-level scenes, it struggles in medium and large-scale scene modeling, producing significant blurring results. Additionally, BAD-GS [52] falls short in regions with significant color and depth variations, and produces overly smooth background textures. Although EDNeRF [3] exhibits overall satisfactory performance, its complex network architecture prolongs the training time (about 7 hours per scene) and precludes real-time rendering. In comparison, our method overcomes the baselines in producing high-fidelity novel views, and significantly reducing training time as well as demonstrating substantial advantages in real-time application scenarios.
Table 2: Quantitative ablation on proposed loss functions. Best-performing results are highlighted in bold and second results in underline.
Large-scale Medium-scale Real Scene
[…]
24.98 .7865 .2986 29.10 .8954 .1673 57.32 8.799 43.26 .1977 6.125
25.71 .7949 .2745 29.94 .9098 .1551 56.16 8.582 42.51 .2009 5.066
25.54 .8003 .2711 30.05 .9139 .1542 56.17 8.499 42.23 .1996 4.923
26.02 .8064 .2680 30.47 .9164 .1519 53.95 8.371 41.52 .2129 4.895
Table 3: Ablation study about the impact of pose optimization.
ATE $\downarrow$ Medium-scale Large-scale
Initial poses $0.0598\pm 0.0167$ $0.3825\pm 0.1245$
Optimized poses $0.0461\pm 0.0129$ $0.0489\pm 0.0112$
Table 4: Robustness against motion blur level.
Blur level Mild Blur Medium Blur Strong Blur
[…]
4.5 Real-world Data Experiments
We present the quantitative results on the captured real-world data in the last row of Table. 1. It can be found that our method achieves superior performance compared to other approaches. Specifically, for NR-IQA metrics, we achieve improvements in BRISQUE [27], NIQE [28], PIQE [40], and RankIQA [21] by 15.38%, 19.50%, 11.49%, and 22.83% respectively. We also achieve an increase in 19.38% in MetaIQA [53]. The qualitative comparisons are shown in the last row of Figure. 3(a) and in the supplementary, which further demonstrate that our method is capable of reconstructing detailed textures, ultimately achieving higher-quality novel view synthesis.
4.6 Ablation Study
@@
Camera Poses Optimization. We firstly conduct ablations to investigate the effect of the number of camera poses optimized in the exposure time. We select five large scenes from our synthetic dataset for evaluation. In the experiments, we vary the number of camera poses, denoted as $n$, from 5, 9, 13, and 17. The quantitative results of the novel view rendering are displayed in Figure. 4. It indicates that the results reach a bottleneck at 9 poses. Beyond this point, the improvements are limited and may potentially lead to local convergence issues. Based on these experiments, we choose $n=9$ camera poses to achieve a balance between rendering performance and training efficiency. Here, we also provide comparison with BAD-NeRF [41] and BAD-GS [52]. These two methods typically use linear interpolation to obtain camera trajectory, while our camera trajectories are estimated from the decomposed latent images, which provides more accurate initialization and helps our method achieves better performance. Moreover, we conduct quantitative experiments using 9 camera poses to compute ATE (Average Trajectory Error) of the initial poses produced by COLMAP [34] and the optimized poses, the results are shown in Table. 3, which validates the effectiveness of pose optimization.
@@
Effectiveness of The Loss Functions. We conduct novel view synthesis experiments on the proposed datasets to validate the effectiveness of the training losses. The quantitative results, as shown in Table. 2, indicate that using only the blur reconstruction loss leads to suboptimal outputs, performing poorly and lacking high-frequency details on both synthetic and real-world datasets. In contrast, incorporating $\mathcal{L}_{\text{event}}$, $\mathcal{L}_{\text{depth}}$, and $\mathcal{L}_{\text{int}}$ enables our proposed method to produce high-fidelity novel views with intricate details.
Robustness Against Motion Blur Levels. To validate the robustness of our method in handling different levels of motion blur, we set up three different camera speeds in the city blocks scene of the synthetic dataset to obtain images with varying degrees of blur. Images with mild blur are captured at half the default camera motion speed, images with medium blur are captured at the default motion speed, and images with strong blur are captured at twice the default motion speed. The quantitative results of novel view synthesis are listed in Table. 4. It can be observed that the results show no significant fluctuations across different levels of blur, demonstrating the robustness of our method to varying motion blur levels. Please refer to the supplementary for more ablations of our method.
Figure 4: Ablation on number of poses in the camera trajectory.